OpenStudy (anonymous):

During one week in the winter, the temperature dropped an average of 2° per hour for 8 hours. If the temperature was -2°F at the beginning of the 8 hours, what was the temperature at the end? A. -20 B. -18 C. -14

OpenStudy (anonymous):

I'm sure I've answered this question here before... \(-2 - 8\times 2 =?\)

OpenStudy (anonymous):

2 per hour, and there were 8 hours, so you do 2 times 8 = 16. it dropped by that amount, so you do -2-16=-18